OpenStudy (anonymous):

Gwen walked 18 miles last week. This week, she walked 4 fewer miles than last week. Write an inequality that compares how far Gwen walked in the two weeks.

OpenStudy (anonymous):

This week is 18 - 4 = 14 miles Gwen walked 18 + 14 = 32 miles

OpenStudy (anonymous):

so it says to write in an inequality is that how it is written?

OpenStudy (anonymous):

I am not sure since it's not really inequality

OpenStudy (lilg132):

\[18 < x \le 32\] dont forget to click good answer

OpenStudy (anonymous):

what

OpenStudy (lilg132):

what?

OpenStudy (anonymous):

you said dont t forget

OpenStudy (lilg132):

i'll rephrase that, remember to click good answer